calculate this quantity in miles per liter : 1.000gal = 3.785 , 1.000km = 0.6214miles; a large suv gets 13.0 miles per gallon of gas

calculate this quantity in kilometers per liter : 1.000gal = 3.785 , 1.000km = 0.6214miles; a large suv gets 13.0 miles per gallon of gas
Answer #1

1)1 gallon = 3.785L

So, 13 miles/gallon = 13 miles/3.785L = 3.434 miles/L (answer)

2)1 km = 0.6214 miles

So, 1 mile = (1/0.6214) km = 1.6092km

and 1 gallon = 3.785L

So, 13miles/gallon = (13*1.6092) km/ 3.785L = 5.526 km/L (Answer)